What are 4'th roots of -8i
Answer by tommyt3rd(5050)

-8i = 8 at 270 deg
n=4
k=0,1,2,3
k=0:
r= 4-th-root(8)at theta=(270)/4=135/2
k=1:
r=4-th-root(8) at theta=(270+360)/4=315/2
k=2:
r=4-th-root(8) at theta=(270+720)/4=495/2
k=3:
r=4-th-root(8) at theta=(270+1080)/4=675/2

:)